A delivery truck arrives at the Robert's store with 8 small boxes and 5 large boxes. The total charge for the boxes, without the tax and delivery charges, is 184 pesos. A large box cost 3 pesos more than a small box. What is the cost of each box?

Answer by nerdybill(7384)

A delivery truck arrives at the Robert's store with 8 small boxes and 5 large boxes. The total charge for the boxes, without the tax and delivery charges, is 184 pesos. A large box cost 3 pesos more than a small box. What is the cost of each box?
.
Let x = cost (pesos) of small box
then
x+3 = cost (pesos) of large box
.
8x + 5(x+3) = 184
8x + 5x+15 = 184
13x+15 = 184
13x = 169
x = 13 pesos (small box)
.
Large box:
x+3 = 13+3 = 16 pesos